### v1

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
The Wineland parameter aims at detecting metrologically useful entangled states, called spin-squeezed states, from expectations and variances of total angular momenta. Efficient strategies for estimating this parameter in practice have yet to be determined. We here formulate the question of whether a state is spin-squeezed according to the Wineland parameter as an hypothesis test and assess the relevance of experimental outcomes by deriving upper and lower bounds on the p-value. By applying our statistical test to data obtained in multiple experiments, we are unable to reject the hypothesis that non-spin squeezed states were measured with a p-value of 5% or less in most cases. We also find an explicit non-spin squeezed state according to the Wineland parameter reproducing most of the observed results with a p-value exceeding 5%. Our results provide a rigorous method to establish robust statistical evidence of spin squeezing from the Wineland parameter in future experiments, accounting for finite-size effects.
Introduction —
Quantum metrology leverages quantum mechanical effects to surpass the limitations of classical measurement techniques [1, 2], notably the standard quantum limit. Among the quantum states that can be utilized to enhance measurement precision [3, 4, 5, 6, 7, 8, 9, 10], spin-squeezed states are receiving significant attention [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The condition for spin-squeezing is usually expressed in terms of the first and second moments of the angular momentum operator which is defined for $N$ distinguishable $1/2$ spins as
$J_{\alpha}=\frac{1}{2}\sum_{i=1}^{N}\sigma_{\alpha}^{(i)}$ (1)
where $\alpha=x,y,z$ specifies the direction and $\sigma_{\alpha}^{(i)}$ is the corresponding Pauli matrix for spin $i$. More precisely, a state that fulfills [23]
$\xi^{2}_{W}=\frac{N\text{ Var}(J_{\vec{n}_{\perp}})}{|\langle J_{\vec{n}}\rangle|^{2}}<1$ (2)
is said to be (metrologically) spin-squeezed along the $\perp$ direction. Here, $\langle J_{\vec{n}}\rangle$ denotes the expectation value of the spin operator along the direction $\vec{n}$ and Var$(J_{\vec{n}_{\perp}})$ is the variance of the spin in a direction $\vec{n}_{\perp}$ orthogonal to $\vec{n}$. $\xi^{2}_{W}$ is called the Wineland parameter. Put simply, $\xi^{2}_{W}<1$ witnesses a class of quantum states where the spin variance is reduced in one direction while being increased in the orthogonal direction [23, 24]. In a Ramsey experiment, $\xi^{2}_{W}$ is established as the ratio of phase sensitivity of the state of interest $\hat{\rho}$ compared to a coherent spin state. Hence $\xi^{2}_{W}<1$ witnesses states having the potential to beat the standard quantum limit. The advantage of spin-squeezed states in metrology is due to quantum correlations between spins [24]. Specifically $\xi^{2}_{W}<1$ witnesses entanglement [25]. The ability to observe multiple spin states that are both metrologically useful and entangled through collective spin measurements has driven numerous experimental implementations with various physical plateforms involving up to almost a million spins [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, the exact values of the variance and expectation of total angular momenta involved in the Wineland parameter cannot be obtained from a finite number of measurement repetitions. Hence, concluding that spin-squeezed states have been created and detected in such systems requires careful consideration due to the limitations of finite statistics. This challenge was highlighted in Refs. [26, 27], which aimed to draw conclusions about Bell correlations in a many-spin system. In this letter, we present a rigorous method to establish robust statistical evidence of spin squeezing from the Wineland parameter, accounting for finite-size effects.
Methods —
To set up the statistical test, the hypotheses need to be formulated. As a first step, we focus on the following linearized form of the Wineland parameter
$\Gamma=N(\Delta Q_{{\vec{n}_{\perp}}})^{2}-\langle Q_{\vec{n}}\rangle^{2}$ (3)
where $Q_{i}=2J_{i}/N$ is the normalized angular momemtum operator in a direction specified by $i$. In the asymptotic limit (infinite number of measurement repetitions), a spin system is considered spin-squeezed if the corresponding state satisfies $\Gamma<0$. Therefore, we define the null hypothesis $H_{0}$ as “the state of interest satisfies $\Gamma\geq 0$”. At first glance, the sample variance and sample mean are natural choices as estimators of $(\Delta Q_{i})^{2}$ and $\langle Q_{i}\rangle^{2}$. However, this results in an estimator for $\Gamma$ that is a nonlinear function of independent random variables, creating challenges in accurately estimating the number of experimental repetitions required to reject the null hypothesis $H_{0}$ without overestimation, see Supplemental Material Sec. I and IV. In this work, we instead introduce the following family of criteria
$\Gamma_{c}=N\langle Q^{2}_{\vec{n}_{\perp}}\rangle-f_{\alpha,\beta}(\langle Q_{\vec{n}_{\perp}}\rangle,\langle Q_{\vec{n}}\rangle).$ (4)
Here, $f_{\alpha,\beta}(x,y)=2\alpha Nx+2\beta y-N\alpha^{2}-\beta^{2}$ is the tangent plane of $h(x,y)=Nx^{2}+y^{2}$ at the point $c=(\alpha,\beta)\in[-1,1]$ and forms an affine lower bound for $h(x,y)$. Therefore, for any given point $(\alpha,\beta)$, $\Gamma_{c}$ serves as a linear upper bound of $\Gamma$, that is, $\Gamma\leq\Gamma_{c}$, and a quantum state satisfying $\Gamma_{c}<0$ is classified as spin-squeezed.
Refer to caption[11][12][12][13][12][14][12][14][12][12][14][15][16][17][18][19][20][21][22]
  • [11] Meyer et al. (2001)
  • [12] Bornet et al. (2023)
  • [13] Franke et al. (2023)
  • [14] Bohnet et al. (2016)
  • [15] Strobel et al. (2014)
  • [16] Riedel et al. (2010)
  • [17] Ockeloen et al. (2013)
  • [18] Schleier-Smith et al. (2010)
  • [19] Leroux et al. (2010)
  • [20] Louchet-Chauvet et al. (2010)
  • [21] Bohnet et al. (2014)
  • [22] Sewell et al. (2012)
Figure 1: Experimental achievements together with their statistical analysis. The experiments that have been reported in the litterature are identified by a square whose position specifies the total number of experimental repetitions ($M$) that has been realized and the system size ($N$). The color of the square gives the observed value of the Wineland parameter. Every square is accompanied by a number used to identify the reference as specified below the graph. The significance of the experimental outcomes are quantified by the number of measurements that are necessary and sufficient to reach a p-value of 5%. Namely, the upper triangle (with the vertex pointing down) gives the total number of measurement repetitions that is sufficient to reject the hypothesis that non-spin squeezed states were measured. The lower triangle (with the vertex pointing up) specifies the number of measurements that is required in order to ensure that a specific non-spin squeeezed state (according to the Wineland parameter) cannot reproduce the observed statistics.
Let $\widetilde{\Gamma}_{{c}}$ be an estimator for $\Gamma_{c}$ and $\gamma_{c}<0$ the observed value in an actual experiment. The significance of the experimental outcome is given by the p-value $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})$. Specifically, If $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})$ is less than a predetermined significance level, such as 0.05, the null hypothesis $H_{0}$ can be rejected and we conclude that the system of interest is spin squeezed.
Establishing upper bounds on the p-value —
For the estimator, we consider the following sum of identical and independent random variables
$\widetilde{\Gamma}_{c}=\sum_{i=1}^{M/2}\Gamma_{c,i}$ (5)
with
$\Gamma_{c,i}=N\left(q^{(i)}_{\vec{n}_{\perp}}\right)^{2}-f_{\alpha,\beta}\left(q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}\right).$ (6)
Here, $q^{(i)}_{\vec{n}_{\perp}}$ $(q^{(i)}_{\vec{n}})$ is the outcome of the measurement operator $2J_{\vec{n}}^{(i)}/N$, $(2J_{\vec{n}_{{\perp}}}^{(i)}/N)$ for experimental round $i$ and $M/2$ is the number of experimental repetitions in each measurement directions ($M$ is the total number of experimental repetitions). To determine the significance level without presuming the probability distribution of the estimator, we employ the Bernstein inequality to bound $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})$, which gives (see Supplemental Material, Sec. II)
$P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})\leq\exp{\dfrac{\gamma^{2}_{c}M/2}{2\Gamma^{c}_{1}\Gamma^{c}_{0}+2(\Gamma^{c}_{1}-\Gamma^{c}_{0})\gamma_{c}/3}}.$ ) . (7)
$\Gamma^{c}_{1}$ and $\Gamma^{c}_{0}$ are the maximal and minimal values of $\Gamma_{c,i}$, which are given by
$\Gamma^{c}_{1}$ $=N(1+|\alpha|)^{2}+(1+|\beta|)^{2}-1,$ (8)
$\Gamma^{c}_{0}$ $=(1-|\beta|)^{2}-1.$
An upper bound on the p-value can thus be obtained in practice by first extracting the value $\gamma_{c}$ that the estimator $\tilde{\Gamma}_{c}$ takes on the observed statistics and then deducing the right hand side of Eq. (7) from the number of experimental repetitions $M$ by fixing $\alpha$ and $\beta$. Ultimately, this upper bound can be tightened by minimizing the right hand side of Eq. (7) over $\alpha$ and $\beta$.
Analysis of existing experimental realizations —
To estimate what could have been achieved in previously implemented experimental realizations, we extracted the value of $\gamma_{c}$ from Refs. [11, 15, 16, 17, 14, 18, 19, 20, 21, 22, 12, 13] by assuming that $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$. We increase $M$, optimizing $(\alpha,\beta)$, until the upper bound on the p-value achieves $5$%. We then execute numerical simulations to show that the assumption $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$ is conservative, i.e. it minimizes the number of experimental repetitions that are needed to achieve a given upper bound on the $p$-value, see Supplemental Material Sec. III. The value of $M$ that we find is reported in Fig. 1 (triangle with the vertex pointing down) as a function of the system size for each experiment separately. The actual number of experimental repetitions conducted is also indicated (square). Except for the experiment involving 16 spins [^1] In Ref. [13], the authors specify that each measurement point reported in their paper is an average of 50 to 600 repetitions of experimental realizations. The position of the corresponding square in Fig. 1 assumes that 1200 repetitions in total were performed, as two measurements directions are involved . Actually, 728 realizations are enough to achieve an upper bound on the p-value of 5%. reported in Ref. [13] and 470 spins [^2] In Ref. [15], a tomography was performed using a total of 32500 experimental realizations. The position of the corresponding square in Fig. 1 hence assumes that 32500 realizations were performed. Actually, 19970 realizations (9985 in each measurement direction) ensure a significance level of 5%. in Ref. [15], we cannot reject the hypothesis that non-spin squeezed states were measured with a p-value less than or equal to 5%. For most experiments, the discrepancy between the actual number of measurements and the number required to achieve an upper bound on the p-value of 5% is at least a factor of 10, and can reach up to $10^{4}$ in the worst case.
Establishing lower bounds on the p-value —
The upper bound on the p-value we derived before suggests that the number of experimental repetitions performed may be insufficient to rule out the possibility that non-spin-squeezed states were actually produced in most existing experimental realisations. To substantiate this claim, we consider the following state
[…]
where |$\ket{\chi}$ is a squeezed state with mean spin direction $\vec{n}$, |$\ket{\pmb{\uparrow_{\vec{n}_{{\perp}}}}}=\ket{\uparrow_{\vec{n}_{{\perp}}}}^{\otimes N}$, |$\ket{\pmb{\downarrow_{\vec{n}_{{\perp}}}}}=\ket{\downarrow_{\vec{n}_{{\perp}}}}^{\otimes N}$ and $0\leq r\leq 1$. The state $\hat{\rho}$ is not considered as being spin squeezed by the Wineland parameter ($\xi_{W}^{2}(\hat{\rho})\geq 1$) for any values of $r$ satisfying
$r\leq r_{\text{max}}=\dfrac{1}{2}\left(\kappa+\sqrt{\kappa^{2}+\dfrac{4N}{\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}}}\right)$ (10)
where $\kappa=\xi_{W}^{2}\left(\ket{\chi}\right)-N/\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}$ (see Supplemental Material Sec. V). When performing $M$ measurements on $\hat{\rho}$ with $r\leq r_{\text{max}}$, the probability to only sample the squeezed state |$\ket{\chi}$ occurs with a probability $p=r^{M}\leq r_{\text{max}}^{M}=p^{*}$. Hence, $p^{*}$ establishes a lower bound on the p-value of a statistical test designed to reject the null hypothesis $H_{0}$.
For each experiment reported in Refs. [11, 15, 16, 17, 14, 18, 19, 20, 21, 22, 12, 13], we compute the minimum number of measurements required to ensure that the observed statistics could not be reproduced by a non-spin-squeezed state of the form given by $\hat{\rho}$ with a probability $p^{*}$ sets to 5%. This is derived from $\log(0.05)/\log(r_{\text{max}})$ where $\xi_{W}^{2}(\ket{\chi})$ and $\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}$ are fixed to the observed values. The results, reported in Fig. 1, show that for most of the experiments involving more than 100 spins, the proposed non-spin squeezed state reproduces the observed results with a probability exceeding 5%.
Discussion and Conclusion —
We designed a linear estimator associated to the Wineland parameter and use the Bernstein inequality to quantify the significance of experimental data. This defines a proper recipe for estimating the Wineland parameter in practice. The first step consists in extracting the value $\gamma_{c}$ that the estimator $\tilde{\Gamma}_{c}$ takes on the observed statistics using Eq. (5) and then deducing the upper bound on the p-value from the minimization over $\alpha$ and $\beta$ of the right hand side of Eq. (7) while fixing the number of experimental repetitions $M$. The null hypothesis associated to “the state of interest satisfies $\Gamma\geq 0$” can then be rejected if the p-value is below 0.05 or any other desired value.
Our analysis also provides an estimate of the number of measurements which is necessary (but not sufficient) to reject the null hypothesis. From a specific state, we showed that the p-value is lower bounded by $p^{*}=r_{\max}^{M}$ where $r_{\max}$ is specified from the expected values of $\xi_{W}^{2}$ and $\langle Q_{\vec{n}}\rangle$, see Eq. (10). Note that $r_{\max}\geq(\sqrt{N^{2}+4N}-N)/2$ and hence, a number of measurements $M\geq\ln{0.05}/{\ln{[(\sqrt{N^{2}+4N}-N)/2]}}\sim-N\log(0.05)$ ) is necessary to achieve a p-value smaller than 5%.
Note that the p-value bounds we proposed could likely be improved. For instance, using more refined estimators may yield tighter bounds. Additionally, various concentration inequalities could be explored, and although several have already been tested (see Supplemental Material), other choices might offer even stronger bounds. Similarly, while we have tested multiple states to construct the lower bound on the p-value, we cannot rule out the possibility that better states exist for this purpose. Moreover, similar methods could be developed to quantify finite-size effects on other squeezing or entanglement parameters [28]. We leave these extensions for future work. Nonetheless, our results already offer a solid framework for predicting and accurately assessing the outcomes of future experiments focused on spin squeezing.
Acknowledgments —
We thank Gregoire Misguich for helpful discussions. We acknowledge funding by the European High-Performance Computing Joint Undertaking (JU) under grant agreement No 101018180 and project name HPCQS and by a French national quantum initiative managed by Agence Nationale de la Recherche in the framework of France 2030 with the reference ANR-22-PETQ-0007 and project name EPiQ.