### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
The Wineland parameter aims at detecting metrologically useful entangled states, called spin-squeezed states, from expectations and variances of total angular momenta. However, efficient strategies for estimating this parameter in practice have yet to be determined and in particular, the effects of a finite number of measurements remain insufficiently addressed. We formulate the detection of spin squeezing as a hypothesis-testing problem, where the null hypothesis assumes that the experimental data can be explained by non-spin-squeezed states. Within this framework, we derive upper and lower bounds on the p-value to quantify the statistical evidence against the null hypothesis. By applying our statistical test to data obtained in multiple experiments, we are unable to reject the hypothesis that non-spin squeezed states were measured with a p-value of 5% or less in most cases. We also find an explicit non-spin squeezed state according to the Wineland parameter reproducing most of the observed results with a p-value exceeding 5%. More generally, our results provide a rigorous method to establish robust statistical evidence of spin squeezing from the Wineland parameter in future experiments, accounting for finite statistics.
Introduction —
Quantum metrology leverages quantum mechanical effects to surpass the limitations of classical measurement techniques [1, 2], notably the standard quantum limit. Among the quantum states that can be utilized to enhance measurement precision [3, 4, 5, 6, 7, 8, 9, 10], spin-squeezed states are receiving significant attention [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The condition for spin-squeezing is usually expressed in terms of the first and second moments of the angular momentum operator which is defined for $N$ distinguishable $1/2$ spins as
$J_{\alpha}=\frac{1}{2}\sum_{i=1}^{N}\sigma_{\alpha}^{(i)}$ (1)
where $\alpha=x,y,z$ specifies the direction and $\sigma_{\alpha}^{(i)}$ is the corresponding Pauli matrix for spin $i$. More precisely, a state that fulfills [23]
$\xi^{2}_{W}=\frac{N\text{ Var}(J_{\vec{n}_{\perp}})}{\langle J_{\vec{n}}\rangle^{2}}<1$ (2)
is said to be (metrologically) spin-squeezed along the $\vec{n}_{\perp}$ direction. Here, $\langle J_{\vec{n}}\rangle$ denotes the expectation value of the spin operator along the direction $\vec{n}$ and Var$(J_{\vec{n}_{\perp}})$ is the variance of the spin in a direction $\vec{n}_{\perp}$ orthogonal to $\vec{n}$. $\xi^{2}_{W}$ is called the Wineland parameter. Put simply, $\xi^{2}_{W}<1$ witnesses a class of quantum states where the collective spin variance is reduced in one direction while being increased in the orthogonal direction [23, 24]. In a Ramsey experiment, $\xi^{2}_{W}$ is established as the ratio of phase sensitivity of the state of interest $\hat{\rho}$ compared to a coherent spin state. Hence $\xi^{2}_{W}<1$ witnesses states having the potential to beat the standard quantum limit. The advantage of spin-squeezed states in metrology is due to quantum correlations between spins [24]. Specifically $\xi^{2}_{W}<1$ witnesses entanglement [25]. The ability to observe multiple spin states that are both metrologically useful and entangled through collective spin measurements has driven numerous experimental implementations involving up to almost a million spins [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
Experimentally confirming spin-squeezing has traditionally relied on the error bar associated to the estimation of the Wineland parameter. However, to demonstrate evidence of spin-squeezing, one should quantify the probability that the observed statistics, obtained from a limited number of experimental repetitions, could result from measurements on a state with no spin-squeezing, i.e. a state satisfying $\xi^{2}_{W}\geq 1$. Error bars only quantify the spread of the measurement outcomes. They do not provide information about the likelihood that the observed data could have been produced by non-spin-squeezed states [26]. A more comprehensive statistical framework is thus needed for testing the Wineland criterion in practice. This challenge was highlighted in Refs. [27], where a similar issue emerged in the context of Bell correlation witnesses in many-spin systems.
In this letter, we develop a rigorous statistical framework for establishing robust evidence of spin squeezing based on the Wineland parameter, explicitly accounting for finite-size effects. Our approach focuses on quantifying the probability (p-value) that the observed experimental data could arise from a non-spin-squeezed state, thereby formulating the problem as a hypothesis test. Unlike the approach in Ref. [28], our analysis does not test a specific non-spin-squeezed state and enables the rejection of the hypothesis that any non-spin squeezed states could have produced the observed statistics with a specified p-value. Our findings indicate that most existing experiments do not sufficiently account for finite-size effects, with only a limited number of cases capable of rejecting the hypothesis test. In many situations, an insufficient number of measurements has been performed which result in a high probability that a non-spin-squeezed state could mimic the experimental outcomes, particularly when the total spin number is large.
Traditional approach —
The verification of spin squeezing in experiments is typically based on the statistical estimation of the Wineland parameter $\xi^{2}_{W}$. Conventional approaches involve measuring the total spin $\langle J_{\vec{n}}\rangle$ and the variance $\text{Var}(J_{\vec{n}_{\perp}})$, followed by an evaluation of the uncertainty in $\xi^{2}_{W}$ from its error bar. A state is conventionally classified as spin-squeezed if the estimated value of $\xi^{2}_{W}$ falls below 1 with a sufficiently small error bar. However, error bars do not constitute a rigorous test for statistical significance [26]. In particular, they fail to account for the probability that statistical fluctuations may cause non-spin-squeezed states to satisfy $\xi^{2}_{W}<1$, leading to potential misclassification. To illustrate this risk, we conduct a Monte Carlo simulation by repeatedly measuring a non-spin-squeezed state of $9$ spins with $\xi_{W}^{2}=1$, as detailed in End Matter. The results show that even with a large number of measurements resulting in a small error bar, statistical fluctuations can still cause a non-spin-squeezed state to be mistakenly classified as spin-squeezed.
Methods —
We here focus on null-hypothesis significance testing, with the p-value quantifying the probability of obtaining a Wineland parameter at least as extreme as the value actually observed, under the null hypothesis $H_{0}$ that the measured state is non-spin-squeezed. A small p-value thus provides strong evidence against $H_{0}$, suggesting that the observed data is unlikely under the assumption that they result from measurements of a non-spin-squeezed state. To ensure the p-value remains independent of any specific quantum state or statistical distribution, we determine an upper bound by considering all possible non-spin-squeezed states.
To formulate the null-hypothesis significance test, we introduce the normalized spin angular momentum operator $Q_{(\cdot)}=2J_{(\cdot)}/N$, define the following parameter
$\Gamma=N\text{Var}(Q_{{\vec{n}_{\perp}}})-\langle Q_{\vec{n}}\rangle^{2}.$ (3)
and use the sample variance and sample mean as unbiased estimators for $\text{Var}(Q_{\vec{n}_{\perp}})$ and $\langle Q_{\vec{n}}\rangle^{2}$. In the asymptotic limit (infinite measurement repetitions), a spin system is considered spin-squeezed if the corresponding state satisfies $\Gamma<0$. While Eq. (3) provides a natural alternative criterion for spin squeezing, it involves nonlinear terms. Existing concentration inequalities struggle to effectively account for these nonlinearities, resulting in an overly conservative upper bound on the p-value. As a result, the predicted number of measurements required to achieve a small p-value is overestimated (see Supplemental Material, Secs. II and V).
To address these challenges, we propose a family of parameters
$\Gamma_{c}=N\langle Q^{2}_{\vec{n}_{\perp}}\rangle-f_{\alpha,\beta}(\langle Q_{\vec{n}_{\perp}}\rangle,\langle Q_{\vec{n}}\rangle),$ (4)
where each element is defined by a fixed value of $c=(\alpha,\beta)$. The function $f_{\alpha,\beta}(x,y)$ is the tangent plan to $Nx^{2}+y^{2}$ at the point $c=(\alpha,\beta)\in[-1,1]$
$f_{\alpha,\beta}(x,y)=2\alpha Nx+2\beta y-N\alpha^{2}-\beta^{2}.$ (5)
Importantly, $f_{\alpha,\beta}(x,y)$ serves as an affine lower bound for $Nx^{2}+y^{2}$ for any $(x,y)$ and $(\alpha,\beta)$. This ensures that $\Gamma_{c}\geq N\langle Q^{2}_{\vec{n}_{\perp}}\rangle-N\langle Q_{\vec{n}_{\perp}}\rangle^{2}-\langle Q_{\vec{n}}\rangle^{2}=\Gamma$ for any choice of $c=(\alpha,\beta)$. As a result, quantum states satisfying $\Gamma_{c}<0$ can be classified as spin-squeezed.
Refer to caption[11][12][12][13][12][14][12][14][12][12][14][15][16][17][18][19][20][21][22]
  • [11] Meyer et al. (2001)
  • [12] Bornet et al. (2023)
  • [13] Franke et al. (2023)
  • [14] Bohnet et al. (2016)
  • [15] Strobel et al. (2014)
  • [16] Riedel et al. (2010)
  • [17] Ockeloen et al. (2013)
  • [18] Schleier-Smith et al. (2010)
  • [19] Leroux et al. (2010)
  • [20] Louchet-Chauvet et al. (2010)
  • [21] Bohnet et al. (2014)
  • [22] Sewell et al. (2012)
Figure 1: Experimental achievements together with their statistical analysis. The experiments that have been reported in the litterature are identified by a square whose position specifies the total number of experimental repetitions ($M$) that has been realized and the system size ($N$). The color of the square gives the observed value of the Wineland parameter. Every square is accompanied by a number used to identify the reference as specified below the graph. The significance of the experimental outcomes are quantified by the number of measurements that are necessary and sufficient to reach a p-value of 5%. Namely, the upper triangle (with the vertex pointing down) gives the total number of measurement repetitions that is sufficient to reject the hypothesis that non-spin squeezed states were measured. The lower triangle (with the vertex pointing up) specifies the number of measurements that is required in order to ensure that a specific non-spin squeezed state (according to the Wineland parameter) cannot reproduce the observed statistics.
To estimate $\Gamma_{c}$ in practice, we consider $M/2$ independent and identically distributed (i.i.d.) repeated processes ($M$ is the total number of experimental repetitions). In each processes $i$, the system undergoes two independent measurements. In the first measurement, the quantum state is prepared, and a measurement of $Q_{\vec{n}}$ is performed, yielding the outcome $q^{(i)}_{\vec{n}}\in[-1,1]$. In the second measurement, the quantum state is re-prepared, and a measurement of $Q_{\vec{n}_{\perp}}$ is performed, obtaining the outcome $q^{(i)}_{\vec{n}_{\perp}}\in[-1,1]$. Using these measurement outcomes, we define the estimator of $\Gamma_{c}$ for each experimental processes $i$ as the random variable
$\widetilde{\Gamma}_{c,i}=N\left(q^{(i)}_{\vec{n}_{\perp}}\right)^{2}-f_{\alpha,\beta}\left(q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}\right).$ (6)
Given the $M/2$ repetitions, $\Gamma_{c}$ can be estimated by using the following statistical estimator
$\widetilde{\Gamma}_{c}=\frac{2}{M}\sum_{i=1}^{M/2}\widetilde{\Gamma}_{c,i}.$ (7)
Let $\gamma_{c}<0$ be the observed value of $\widetilde{\Gamma}_{c}$ from a given experimental realization. We are interested in the p-value, i.e. the probability that the estimator $\widetilde{\Gamma}_{c}$ takes a value equal to or smaller than $\gamma_{c}$, assuming the measurements are performed on non-spin-squeezed states (the null hypothesis $H_{0}$). We name it explicitly as $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$. Importantly, $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$ must be estimated by considering all possible non-spin-squeezed states and all possible statistical distributions. If $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$ is less than a predetermined significance level, such as 0.05, we conclude that the system of interest exhibits spin squeezing.
Establishing upper bounds on the p-value —
Concentration inequalities [29, 30] can be used to upper bound $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$. As shown in Supplemental Material, Sec. III, the Bernstein inequality for example gives
$P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})\leq\exp{\dfrac{\gamma^{2}_{c}M/2}{2\Gamma^{c}_{1}\Gamma^{c}_{0}+2(\Gamma^{c}_{1}-\Gamma^{c}_{0})\gamma_{c}/3}}$ ) (8)
where
$\Gamma^{c}_{1}$ $=\max_{q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}}\widetilde{\Gamma}_{c,i}=N(1+|\alpha|)^{2}+(1+|\beta|)^{2}-1,$
$\Gamma^{c}_{0}$ $=\min_{q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}}\widetilde{\Gamma}_{c,i}=(1-|\beta|)^{2}-1.$
The right-hand side of Ineq. (8) is solely a function of the parameter $c=(\alpha,\beta)$ and the experimental observation $\gamma_{c}$. It can be minimized over $c=(\alpha,\beta)$ under the constraint that $\gamma_{c}<0$ in order to get the tightest upper bound on the p-value. This defines $p_{\text{opt}}$ as
$p_{\text{opt}}=\inf_{\alpha,\beta\text{ s.t.}\gamma_{c}<0}\exp{\dfrac{\gamma^{2}_{c}M/2}{2\Gamma^{c}_{1}\Gamma^{c}_{0}+2(\Gamma^{c}_{1}-\Gamma^{c}_{0})\gamma_{c}/3}}.$ ) . (9)
Guidelines for robust statistical evidence of spin squeezing in future experiments —
We now have all the ingredients to propose a concrete recipe that can be used to establish a robust statistical evidence of spin squeezing in future experiments. First, we consider $M/2$ measurements of both $Q_{\vec{n}}$ and $Q_{\vec{n}_{\perp}}$. This yields outcomes $q^{(i)}_{\vec{n}}$ and $q^{(i)}_{\vec{n}_{\perp}}$ for $i\in[1,M/2]$. $\gamma_{c}$ is then deduced from the estimator $\widetilde{\Gamma}_{c}$ given in Eq. (6). An upper bound on the p-value is obtained from the minimization given in Eq. (9) over $\alpha$ and $\beta$ such that $\gamma_{c}<0$. A small $p_{\text{opt}}$ provides strong evidence against $H_{0}$, suggesting that the observed data is unlikely under the assumption that they result from measurements of a non-spin-squeezed state.
Analysis of existing experimental realizations —
To estimate what could have been achieved in previously implemented experimental realizations, we calculated the value of $\gamma_{c}$ from Refs. [11, 15, 16, 17, 14, 18, 19, 20, 21, 22, 12, 13] by assuming that $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$. We increase $M$, optimizing $(\alpha,\beta)$, until the upper bound on the p-value achieves $5$%. We then execute numerical simulations to show that the assumption $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$ is conservative, i.e. it minimizes the number of experimental repetitions that are needed to achieve a given upper bound on the $p$-value, see Supplemental Material Sec. IV. The value of $M$ that we find is reported in Fig. 1 (triangle with the vertex pointing down) as a function of the system size for each experiment separately. The actual number of experimental repetitions conducted is also indicated (square). Except for the experiment involving 16 spins [^1] In Ref. [13], the authors specify that each measurement point reported in their paper is an average of 50 to 600 repetitions of experimental realizations. The position of the corresponding square in Fig. 1 assumes that 1200 repetitions in total were performed, as two measurements directions are involved . Actually, 728 realizations are enough to achieve an upper bound on the p-value of 5%. reported in Ref. [13] and 470 spins [^2] In Ref. [15], a tomography was performed using a total of 32500 experimental realizations. The position of the corresponding square in Fig. 1 hence assumes that 32500 realizations were performed. Actually, 19970 realizations (9985 in each measurement direction) ensure a significance level of 5%. in Ref. [15], we cannot reject the hypothesis that non-spin squeezed states were measured with a p-value less than or equal to 5%. For most experiments, the discrepancy between the actual number of measurements and the number required to achieve an upper bound on the p-value of 5% is at least a factor of 10, and can reach up to $10^{4}$ at the worst case.
Establishing lower bounds on the p-value —
The upper bound on the p-value we derived before suggests that the number of experimental repetitions performed may be insufficient to rule out the possibility that non-spin-squeezed states were actually produced in most existing experimental realisations. To substantiate this claim, we follow the approach of Refs. [31, 28] and in particular, we consider the following state
[…]
where |$\ket{\chi}$ is a squeezed state with mean spin direction $\vec{n}$, |$\ket{\pmb{\uparrow_{\vec{n}_{{\perp}}}}}=\ket{\uparrow_{\vec{n}_{{\perp}}}}^{\otimes N}$, |$\ket{\pmb{\downarrow_{\vec{n}_{{\perp}}}}}=\ket{\downarrow_{\vec{n}_{{\perp}}}}^{\otimes N}$ and $0\leq r\leq 1$. The state $\hat{\rho}$ is not considered as being spin squeezed by the Wineland parameter ($\xi_{W}^{2}(\hat{\rho})\geq 1$) for any values of $r$ satisfying
$r\leq r_{\text{max}}=\dfrac{1}{2}\left(\kappa+\sqrt{\kappa^{2}+\dfrac{4N}{\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}}}\right)$ (11)
where $\kappa=\xi_{W}^{2}\left(\ket{\chi}\right)-N/\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}$ (see Supplemental Material Sec. VI). When performing $M$ measurements on $\hat{\rho}$ with $r\leq r_{\text{max}}$, the probability to only sample the squeezed state |$\ket{\chi}$ occurs with a probability $p=r^{M}\leq r_{\text{max}}^{M}=p^{*}$. Hence, $p^{*}$ establishes a lower bound on the p-value of a statistical test designed to reject the hypothesis that the measured state is of the form given in Eq. (10[31, 28].
For each experiment reported in Refs. [11, 15, 16, 17, 14, 18, 19, 20, 21, 22, 12, 13], we compute the minimum number of measurements required to ensure that the observed statistics could not be reproduced by a non-spin-squeezed state of the form given by $\hat{\rho}$ with a probability $p^{*}$ sets to 5%. This is derived from $\log(0.05)/\log(r_{\text{max}})$ where $\xi_{W}^{2}(\ket{\chi})$ and $\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}$ are fixed to the observed values. The results, reported in Fig. 1, show that for most of the experiments involving more than 100 spins, the proposed non-spin squeezed state reproduces the observed results with a probability exceeding 5%.
Discussion and Conclusion —
In this paper, we developed a rigorous analytical framework for detecting spin-squeezed states, with a particular focus on the impact of finite measurement repetitions on the detection results. Specifically, by linearizing $\xi_{W}^{2}$, we proposed a family of criteria for identifying spin-squeezed states. Using a hypothesis-testing approach, we derived upper bounds on the p-value, which can be used to provide a reliable statistical evidence of spin squeezing.
The effects of finite measurement repetitions have long been overlooked in spin-squeezing detection experiments. Our hypothesis-testing framework reveals that most existing experiments fail to pass the statistical test due to an insufficient number of measurements. This issue becomes particularly severe when the spin number is large, as the required number of measurements needed to reliably reject the null hypothesis increases significantly, often exceeding the practical capabilities of current experimental setups. However, this does not necessarily mean that spin squeezing is absent, but rather that the available data does not provide strong enough statistical support for claiming spin squeezing is produced.
Our analysis provides an estimate of the number of measurements which is necessary (but not sufficient) to reject the null hypothesis. From a specific state, we showed that the p-value is lower bounded by $p^{*}=r_{\max}^{M}$ where $r_{\max}$ is specified from the expected values of $\xi_{W}^{2}$ and $\langle Q_{\vec{n}}\rangle$, see Eq. (11). Note that $r_{\max}\geq(\sqrt{N^{2}+4N}-N)/2$ and hence, a number of measurements $M\geq\ln{0.05}/{\ln{[(\sqrt{N^{2}+4N}-N)/2]}}\sim-N\log(0.05)$ ) is necessary to achieve a p-value smaller than 5%.
Note that the p-value bounds we proposed in Eq. (9) could likely be improved. For instance, using more refined estimators may yield tighter bounds. Additionally, various concentration inequalities could be explored, and although several have already been tested (see Supplemental Material Sec. V), other choices might offer even stronger bounds. Similarly, while we have tested multiple states to construct the lower bound on the p-value, we cannot rule out the possibility that better states exist for this purpose. Moreover, our statistical method could be extended to quantify finite-size effects on other squeezing or entanglement criteria [32]. We leave these extensions for future work. Nonetheless, our results already offer a solid framework for predicting and accurately assessing the outcomes of future experiments focusing on spin squeezing.
Acknowledgments —
We thank Gregoire Misguich for helpful discussions. We acknowledge funding by the European High-Performance Computing Joint Undertaking (JU) under grant agreement No 101018180 and project name HPCQS and by a French national quantum initiative managed by Agence Nationale de la Recherche in the framework of France 2030 with the reference ANR-22-PETQ-0007 and project name EPiQ.
Appendix A End Matter
In experimental estimations of the Wineland parameter, standard error bars reflect only the statistical variability of the measured data and do not rigorously quantify the likelihood that the observed statistics could originate from non-spin-squeezed states. To highlight this limitation, we performed Monte Carlo simulations on a specific non-spin-squeezed state, demonstrating explicitly that small error bars do not necessarily correspond to small p-values. This reveals the potential risk of misclassifying non-spin-squeezed states as spin-squeezed when relying solely on error bar.
Specifically, we consider quantum states of the form given by Eq. (10)
[…]
where $|\chi\rangle$ is generated via an appropriate time evolution under a dipolar XY Hamiltonian (see Supplemental Material Sec. I). In our simulations, we use a square lattice of $N=9$ spins and obtain a spin-squeezed state characterized by $\xi_{W}^{2}(|\chi\rangle)\approx 0.41036$, a mean spin polarization $\langle 2J_{y}/N\rangle_{\chi}\approx 0.80812$, and a variance $\text{Var}(2J_{x}/N)_{\chi}\approx 0.02977$. From the value of $\xi_{W}^{2}(|\chi\rangle)$, we construct a non-spin-squeezed state $\hat{\rho}$ satisfying $\xi_{W}^{2}(\hat{\rho})=1$, corresponding to parameters $r\approx 0.96154$, $\langle 2J_{y}/N\rangle_{\rho}\approx 0.77704$ and $\text{Var}\left(2J_{x}/N\right)_{\rho}\approx 0.06709$.
The statistics that would be obtained from such a state are generated using Monte Carlo sampling. Two data sets, each of size $M/2$, are produced ($M$ is the total number of samples). The first data set, $Q_{y}=\{y_{1},\dots,y_{M/2}\}$, is sampled in the eigenbasis of $2J_{y}/N$. The second data set, $Q_{x}=\{x_{1},\dots,x_{M/2}\}$, is sampled in the eigenbasis of $2J_{x}/N$. To establish a comprehensive error estimation framework, the measurement outcomes $x_{i}$ and $y_{j}$ are assumed to be independent and identically distributed (i.i.d.). Their (unknown) true means are labeled ($\mu_{x}$, $\mu_{y}$) and true variances by ($\sigma_{x}^{2}$, $\sigma_{y}^{2}$).
Recalling the definition of the Wineland parameter from Eq. (2), it can be conveniently rewritten as
$\xi_{W}^{2}=\frac{N\left(\mu_{x^{2}}-\mu_{x}^{2}\right)}{\mu_{y}^{2}},$ (13)
where $\mu_{x^{2}}$ is the true mean of the second-order moment of $x_{i}$ which satisfies $\mu_{x^{2}}=\sigma_{x}^{2}+\mu_{x}^{2}$. A natural approach to estimate $\xi_{W}^{2}$ in practice is to substitute the true expectations by the sample means. However, this introduces a systematic bias due to the nonlinear dependence on $\mu_{x}$ and $\mu_{y}$ [33]. To mitigate this issue, the jackknife method can be used [13, 15]. Specifically, the bias-corrected jackknife estimation of $\xi_{W}^{2}$ is given by (see Supplemental Material Sec I)
$\frac{M}{2}\xi_{W}^{2}(\bar{x},\overline{x^{2}},\bar{y})-\left(\frac{M}{2}-1\right)\overline{[\xi_{W}^{2}]},$ (14)
where $\overline{x}=\frac{1}{M/2}\sum_{i=1}^{M/2}x_{i}$, $\overline{x^{2}}=\frac{1}{M/2}\sum_{i=1}^{M/2}x_{i}^{2},$ $\overline{y}=\frac{1}{M/2}\sum_{i=1}^{M/2}y_{i}$ are the sample means of datasets. $\overline{[\xi_{W}^{2}]}$ is the mean of jackknife replicates, defined as $\overline{[\xi_{W}^{2}]}=\frac{1}{M/2}\sum_{i=1}^{M/2}\xi_{W}^{2}\left([x]_{i},[x^{2}]_{i},[y]_{i}\right)$ with $[x]_{i}=\frac{1}{M/2-1}\sum_{j\neq i}x_{j}$, $[y]_{i}=\frac{1}{M/2-1}\sum_{j\neq i}y_{j}$ the jackknife leave-one-out averages.
The uncertainty in the jackknife estimation is quantified by the error bar
$\sqrt{M/2-1}\times\sqrt{\overline{[\xi_{W}^{2}]^{2}}-\left(\overline{[\xi_{W}^{2}]}\right)^{2}},$ (15)
where the second moment of the jackknife replicates is given by $\overline{[\xi_{W}^{2}]^{2}}=\frac{1}{M/2}\sum_{i=1}^{M/2}\left[\xi_{W}^{2}\left([x]_{i},[x^{2}]_{i},[y]_{i}\right)\right]^{2}$. This error bar scales as $O(1/\sqrt{M/2})$, substantially exceeding the bias in the estimate (see Supplemental Material Sec. I).
Figure 2: (Top) Estimation of the Wineland parameter from Monte Carlo samples of size $M$ sampled from a non-spin-squeezed state. Two data sets were sampled: one in the eigenbasis of the mean spin direction ($J_{y}$) and another in the eigenbasis of the squeezing axis. Each error bar corresponds to one standard deviation. (Bottom) P-value upper bound using the method proposed in the main text on the Monte-Carlo samples.
Figure 2 shows the results of the Monte Carlo simulation with the jackknife estimation and uncertainty with one standard error, i.e.
$\widetilde{\xi^{2}_{W}}$ $=(M/2)\xi_{W}^{2}\left(\bar{x},\overline{x^{2}},\bar{y}\right)-(M/2-1)\overline{[\xi_{W}^{2}]}$ (16)
±$\pm\sqrt{M/2-1}\times\sqrt{\overline{[\xi_{W}^{2}]^{2}}-\left(\overline{\xi_{W}^{2}}\right)^{2}}$
as a function of the total number of samples $M$. Although these values fluctuate around the true value $\xi^{2}_{W}=1$, some are noticeably below 1. For instance, for $M=400$, we observe $\widetilde{\xi^{2}_{W}}=0.65832\pm 0.13115$, despite the data being generated by a non-spin-squeezed state. The same figure also displays the p-value as a function of the number of measurements. None of the p-value drops below 0.25037. In particular, the p-value for $M=400$ is below 0.66661 clearly indicating a high probability that the measured state is non-spin-squeezed. More generally, none of these results can pass the hypothesis test and thus do not provide statistically significant evidence for spin squeezing. These simulation results (see also the complementary results presented in Supplemental Material Sec. I) emphasize that “error bars primarily quantify the spread of the data or the precision of the mean estimate, but do not constitute a rigorous test for statistical significance [26]”. Therefore, to provide a more reliable assessment of spin squeezing, it is essential to incorporate p-values as a complementary statistical measure to properly evaluate the probability that a non-spin squeezed state reproduces the observed statistics.